### v3

# Survey in Characterization of Semantic Change

Abstract
Live languages continuously evolve to integrate the cultural change of human societies. This evolution manifests through neologisms (new words) or semantic changes of words (new meaning to existing words). Understanding the meaning of words is vital for interpreting texts coming from different cultures (regionalism or slang), domains (e.g., technical terms), or periods. In computer science, these words are relevant to computational linguistics algorithms such as translation, information retrieval, question answering, etc. Semantic changes can potentially impact the quality of the outcomes of these algorithms. Therefore, it is important to understand and characterize these changes formally. The study of this impact is a recent problem that has attracted the attention of the computational linguistics community. Several approaches propose methods to detect semantic changes with good precision, but more effort is needed to characterize how the meaning of words changes and to reason about how to reduce the impact of semantic change. This survey provides an understandable overview of existing approaches to the characterization of semantic changes and also formally defines three classes of characterizations: if the meaning of a word becomes more general or narrow (change in dimension) if the word is used in a more pejorative or positive/ameliorated sense (change in orientation), and if there is a trend to use the word in a, for instance, metaphoric or metonymic context (change in relation). We summarized the main aspects of the selected publications in a table and discussed the needs and trends in the research activities on semantic change characterization.
1 Introduction
Language seems to be one of the most dynamic skills that humans have acquired through their evolution. It has constantly evolved to adapt the communication to new purposes and needs (Pinker,, 2003, p. 13-23). While there is no single unified theory about its origins (Allan,, 2013, p. 13-52), it is clear that language has undergone significant modifications since the early days of human history. One of the prevailing theories of language creation and evolution is based on the assumption of “continuity of change” (Rilling et al.,, 2012). This theory holds that language is too complex to have arisen suddenly, and must have gradually evolved from earlier pre-linguistic systems used by our ancestors (Pinker and Bloom,, 1990, p. 725-726). In this context, comprehending the process of language change is an contribution to the fields of history, sociology, linguistics, philosophy, and Natural Language Processing (NLP) (Campbell,, 2013, p. 1-8), as it is essential to understanding the origins of language, cultural development, grammatical theories, and the relation between language and thought (Sherwood et al.,, 2008, p. 426–454).
Many unplanned factors drive language change. For example, linguistic advances, social, cultural, and historical influences, as well as technical improvements. There is a variety of grammatical factors that also promote changes, including:
  • Phonological Changes changes in the creation, perception, and representation of sounds in language. For instance, the pronunciation of vowel sounds significantly changed due to the Great Vowel Shift in Medieval English (Nevalainen and Traugott,, 2016, p. 367).
  • Orthographic Changes alterations in the written form of words to form meaningful sentences and to pair with phonological changes, e.g., panem to pain, ‘bread’ in Latin to French.
  • Syntactic Changes alterations to the arrangement of words to create coherent sentences. For example, Modern English emphasizes word order and auxiliary verbs more than the intricate inflectional system used in Old English to identify grammatical connections (Nevalainen and Traugott,, 2016, p. 616).
  • Semantic Changes phrases or words that have meaning changed over time, e.g., notebook “a book for notes”, to “a portable computer.”
While alterations in sound and form are crucial elements shaping our language, semantics plays a pivotal role in driving these changes. Semantic drift influences how words adapt to evolving cultural, technological, and societal contexts, creating new meanings, and preserving the language’s fundamental purpose of effective communication. Despite being a recurring topic in scientific research, semantic drifts remain less understood than other forms of linguistic evolution due to their often subtle and less observable nature, such as those influenced by grammar reforms (Allan and Robinson,, 2012, p. 2-3).
Historically, their approach to studying this process primarily relied on manual methods. Linguists would painstakingly analyze historical texts, dictionaries, and linguistic corpora, tracing the shifting meanings of words through centuries of written records. These manual studies provided invaluable insights into the intricacies of semantic change but were labor-intensive, time-consuming, and difficult to reproduce.
However, with the advent of computational linguistics and the development of advanced computer models, the field underwent a transformative shift. Linguists began harnessing the power of machine learning and natural language processing to automate the analysis of vast amounts of textual data. So the research community gradually shifted from anthropological observation to statistical methods for language analysis (Allan and Robinson,, 2012, p. 161-164).
These computer models, fueled by large corpora, and sophisticated algorithms, enabled researchers to detect subtle properties in language with unprecedented accuracy and efficiency. Enabling studying languages without any previous knowledge of it’s syntax or semantics, comparing multiple languages between themselves in the pursuit for universal properties (Vulic et al.,, 2020).
By embracing a model-based approach, linguistics can now unravel the intricacies of language phenomena and cope with the complexity of linguistics from different cultures and backgrounds (Nivre et al.,, 2020). This scientific tooling provides a clear pathway for progress by facilitating direct comparisons between hypothesis, mathematical models, their refinement, and testing against emerging linguistic phenomena (Hammarström et al.,, 2019).
While initially only observed with a theoretical motivation, the practical importance in computer science became very relevant. Historical knowledge graphs can contain thousands of entities and relations that change over time (e.g., Wikidata and English WordNet) – automatically identifying concepts that change can allow an automatic update of these systems and enable semantic resources to extend the taxonomy of preexisting concepts (Polleres et al.,, 2023).
Other examples of applications that are impacted by semantic change:
  • Personalized vernacular voice assistants. In Brazil, ‘bicha’ can mean “homosexual,” “unidentified female,” or “queue,” depending on which state this assistant is located in. Therefore, it is desirable for an agent using vernacular communication to be aware of this distinction.
  • Region-based translation. In French, ‘déjeuner’ can mean “breakfast,” “lunch,” or “dinner,” depending on which French-speaking country you are addressing the translation. So, a correct translation should respect the geographic origin of the text.
  • Social network analysis. Some political labels have different meanings and orientations depending on which sub-network you analyze (e.g., “We need to change. MAGA!” meaning the political slogan, with a positive sense or “He is a MAGA” meaning he support a politician who uses this slogan, with a negative sense). So, under-specified sentences could mislead sentiment analysis systems.
  • Merging domain ontologies. The word ‘machine’ can mean “a computer program” in computer science literature or “a mechanical device” in engineering literature. So, the entity ‘machine’ does not hold all the relations between both ontologies.
Addressing this issue involves continuously updating NLP models and algorithms to account for semantic change, incorporating real-time data sources, and staying aware of evolving language trends (Lu et al.,, 2022). It underscores the need for adaptability, ongoing maintenance and engineering, and explain decisions of NLP applications that rely on natural language understanding. Specially in the case of Language Models where regular updates requires substantial computational resources, time, and financial investment.
The development of cost-effective tools that can handle the many variations in language and precisely understand the change in semantics is crucial for many businesses. However, while the linguistic field has broadly studied what phenomena cause change (Traugott,, 2017), and many works provide deep analysis and theories on types of change. The computational linguistics community performed considerable effort to detect these drifts, nonetheless, to the best of our knowledge, less effort has been made to formalize and understand what kind of drift a word has undergone (Hengchen et al.,, 2021).
Another important development in this field is a proper language and formalism to describe the problems in semantic change (Kutuzov et al.,, 2018). Having a formal language (i.e., a mathematical description) is of paramount importance in a scientific field for several reasons. It provides a structured framework that allows researchers to precisely articulate and communicate their hypotheses and findings. It also ensures that ideas are expressed in a clear and unambiguous manner, facilitating collaboration and knowledge dissemination within the field. Furthermore, employing a mathematical language can significantly enhance computational interpretation of semantic change, advancing our comprehension of one of the most intriguing aspects of human communication.
This survey addresses the need to understand and formalize semantic change’s characterization. Characterizing such changes will open new perspectives in various areas, mainly because we have no resources to study them [^1] While NLP resources have hundreds of thousands of examples, the datasets of semantic change generally cover less than a hundred words Kutuzov et al., (2022).. This is especially problematic in contexts where both word meanings are relevant, potentially leading to political or social misinterpretations. The survey’s objectives include:
  1. 1.
    Identifying the limitations of current representations and methodologies in the categorization of semantic changes.
  2. 2.
    Analyzing the definitions of each category and the complexity of computer algorithms’ interpretations.
  3. 3.
    Evaluating the strengths of approaches used to classify lexical semantic changes (LSC).
  4. 4.
    Finding and analyzing formal definitions based on first-order logic to reduce ambiguity and facilitate the comparison of findings.
There are many distinct categories for semantic change, ranging from cultural tradition (e.g., taboo) to cognitive aspects (e.g., metaphor). On this survey we select categories that are both addressable by computational methods on a comparative setting (Tang et al.,, 2013) and by linguistic literature (Campbell,, 2013, p. 221-246), i.e., changes we can systematically measure and are backed by a theoretical analysis (Hammarström et al.,, 2019). We will focus on three factors of semantic changes: the orientation (i.e., if the words meaning is getting more pejorative or positive), the relation (i.e., if the words meaning change through a figurative usage), and the dimension (i.e., if the words meaning is becoming more abstract or more specific). Some of these properties, as we detail later, are crucial, for instance, for understanding and explaining the intention of a sentence, for writing tools to assist translation or to interact with users (e.g. in social robot). Finally, it will open new opportunities to improve reasoning capabilities for NLP tools (Danilevsky et al.,, 2020).
This survey is the first to comprehensively cover semantic change characterization by consolidating and categorizing existing approaches and identifying gaps. It introduces formal descriptions and proposes enhancements for a more robust understanding of semantic change. By synthesizing prior research and addressing its limitations, the survey streamlines existing work and advances our understanding of this complex phenomenon.
The remainder of the paper is structured as follows: Section 2 reviews existing surveys and their difference with respect to this work. In Section 3 we investigate available corpora of documents for semantic change identification, characterization and their limitations. In Section 4 we explore existing methods to represent meaning change characterization and how they attempt to solve the characterization problem. Section 5 we present a formal model derived from reviewed methodologies and linguistic description. Section 6 presents our analysis and open questions in characterization. Finally, Section 7 we conclude this survey and outline future work.
2 Lexical Semantic Change
Concerning the characterization of change, we adopt in this survey the mainstream typology in linguistic literature for semantic change (Traugott,, 2017), (i) broadening and narrowing, (ii) amelioration and pejoration, and (iii) metonymization and metaphorization. Before starting our study on change characterization, we reviewed the works published on change identification. This is a well-developed topic that focuses on evaluating whether the meaning of a word has changed or not. Papers addressing change identification often claim the need for further understand and characterization of the type of change that occurred, but the current state of works on computational change characterization is unclear. This conclusion motivated the work presented in this paper and differentiate our survey from existing ones on Lexical Semantic Change (LSC).
In the following subsections we present existing surveys dealing with the semantic change problem and identify the gaps of the current literature. Inspired by the definitions introduced in these studies, we propose our own typology of changes. The last section introduces the objectives of this survey and how we plan to address the gap in semantic change characterization in our future work.
2.1 Related Works on Identification of Lexical Semantic Changes
Surveys for semantic change identification are regularly published, with a main focus on methodologies for change detection. In a survey about “semantic change computation” (Tang,, 2018), Tang proposes to split LSC analysis into 4 steps: obtaining and preparing a diachronic corpus, identifying word senses, modeling semantic change, followed by presentation and validation of hypotheses. This framework sets the structure of the survey, describing the studies related to each of these steps. The characterization of changes is mainly discussed in the “modeling semantic change” section, where the author explains the difference between Cambridge Change (Ardley,, 1968, p. 91), and McTaggart Change (Broad,, 1938, p. 297) and cites the approaches following these definitions. However, the approaches that can classify semantic changes as sense broadening, sense narrowing, pejoration, amelioration, metaphorical change and metonymical change are not included.
The survey about computational approaches to LSC (Tahmasebi et al.,, 2018) focuses on word-level differentiation, sense differentiation, and lexical replacement. While it includes approaches with several types of change categories (e.g., metaphor, metonymy), the main focus of change categories are broadening, narrowing and correlated types (split, join, birth). However, one recommendation of the authors is to have deepen studies on methods to automatically detecting other types of change and to use it to build the history of the words. Auxiliary to the methodologies discussed, the authors provide an in-depth analysis of linguistic literature that grounds each domain, followed by evaluation strategies and applications. The difficulty to compare approaches is highlighted in the survey and, according to the authors, is the result of an absence of standard metrics/practices and reference datasets.
Kutuzov et al., (2018) survey word representation models and data for semantic change identification, and describe a framework to compare these methodologies. The authors focus on defining a consistent terminology, and coherent practices to evaluate these models. In this work, they claim that there’s still a gap in studying sub-classifications of change:
“However, more detailed analysis of the nature of the shift is needed. […] This problem was to some degree addressed by Mitra et al., (2015), but much more work is certainly required to empirically test classification schemes proposed in much of the theoretical work […]”
First studies on the field employed graph- or frequency-based methods to analyse LSC, later approaches benefited from the advances in the state-of-the-art to replace the analytical methods by those based on embeddings and contextual embeddings. Although contextual embeddings represented a huge improvement for a series of NLP tasks (Devlin et al.,, 2019), the improvement for LSC did not follow this trend (Schlechtweg et al.,, 2020), showing that there is still room for improvement in this domain (Zhou et al.,, 2021; Kutuzov et al.,, 2022).
In this context, Dubossarsky et al., (2017) investigate the quality of laws addressing the properties of change. The authors observed that laws proposed in previous studies are affected by the bias in the model or in the data. When this bias was reduced, the ‘laws’ either did not hold or were less precise than originally reported.
Research in this field also delves into ongoing LSC methodological issues. These issues involve the challenges of characterizing semantic shifts. As pointed out by Hengchen et al., (2021), there are few studies that detail the nature of these changes, mainly because it is difficult to model meaning from textual data. It further concludes that no particular measures for quantifying category change exist and that a consensus on taxonomy for classifying types of change is required.
In the work of Montanelli and Periti, (2023), the authors review methodologies based on contextualized word embeddings. They categorize methodologies for training contextualized embeddings based on how they represent meaning, how the time information is presented to the learning model, and what kind of learning procedure is performed (unsupervised or supervised). In this paper, the authors also discuss the importance of characterization:
“Most of the literature papers do not investigate the nature of the detected shifts, meaning that they do not classify the semantic shifts according to the existing linguistic theory (e.g., amelioration, pejoration, broadening, narrowing, metaphorisation, and metonymisation) […]. These studies could be crucial to detect “laws” of semantic shift that describe the condition under which the meanings of words are prone to change. […]”
For this survey, we argue that semantic change characterization needs to have appropriated methods and tools to analyse each fine-grained category. We also defend that observing variations in a single property can not describe well all forms of LSC. For instance, changes in angles between vectors representing words (Dubossarsky et al.,, 2017) is not the best solution to identify narrowing/broadening of sense (Tahmasebi et al.,, 2018). Specific methods for identifying the different types of change could lead to a better understanding of LSC than straightforward dissimilarity measures.
This survey addresses the need to depict the characterization problem in semantic change and present recent works dealing with this problem. In the next section we define the survey objectives and search criteria to select papers for change characterization.
2.2 Objectives
Different terms are used to refer to Lexical Semantic Change (LSC), such as semantic drift or semantic shift. In this paper we make the distinction between identification of a change and characterization of the change. The former refers to the process through which the meaning of a word gradually or significantly changes over time in a language. While the following refers to which characteristic could describe and differentiate the process that lead to a change in meaning. We formally differentiate this setting in Section 5.1. The analysis of works dealing with the identification problem has been well covered by Tahmasebi et al., (2018), Kutuzov et al., (2018), and Tang, (2018).
Semantic change occurs by the need to reduce fuzziness when combining meanings and forms (Hock and Joseph,, 2019). While some authors propose to use the cause of the changes to group them, other authors categorize the type of change by language figures that describe the change, or by the cultural process. The focus of our literature review is on the characterization of change, where we look for common properties that allows for grouping identified semantic changes into categories. From a computation perspective, determining the cause of a semantic change becomes very complex to implement and, to the best of our knowledge, none computational approach was implemented yet.
Some frameworks were proposed on how to pose and track LSC like Cambridge and McTaggart (Tang et al.,, 2013), which initially were not considering computational methods to it. Cambridge (Ardley,, 1968) and McTaggart (Broad,, 1938) change approaches are different on how to use temporal information to identify changes, while Cambridge proposes to compare two different corpus (from different time period), McTaggart prefers to study the evolution of the usage of a word over a period of time. The approach that best suits the computational perspective is the Cambridge, as McTaggart approach would require historiographer work to acquire historic socio-cultural knowledge that goes beyond contextual analysis of word occurrence (Tang et al.,, 2013, p. 4-7).
For this review we assume the Cambridge setting for change analysis using language figures as bases to create categories of changes, which align with current state-of-the-art methods for computational LSC (Pivovarova and Kutuzov,, 2021). Given this considerations, we adopt the Traugott, (2017) typology, which is also the most predominant categories in literature, as in Koch, (2016, p. 21-66), Hock and Joseph, (2019, p. 189-222), and Campbell, (2013, p. 221-246):
  • Broadening gaining a new meaning, that is not related or not, to the previous meaning, such that a symbol represents more concepts, e.g., ‘cloud’ a computing infrastructure.
  • Narrowing restriction of meaning where a symbol represents fewer concepts, as in ‘gay’ that used to have senses as festive or happy, now it generally is used in a single sense of homosexual.
  • Amelioration a word gained a more positive sense with respect to the previous sense, nice, ‘foolish, innocent’ changed to ‘pleasant.’
  • Pejoration the word is used with a worse connotation with respect to the previous usage, stincan, ‘smell (sweet or bad)’ changed to stink.
  • Metonymization association between terms, e.g., board ‘table’, changed to “people sitting around a table, governing body.”
  • Metaphorization conceptualizing one thing in terms of another, e.g., the use of Latin ad ’to’ + mit ’send’ for locution (admit) (Traugott,, 2017).
As the types in the above typology are directly correlated (involves the same computational challenge), e.g., broadening-narrowing, requires sense counting, amelioration-pejoration, requires sentiment analysis, metaphorization-metonymization, requires some form of relatedness measure (Koch,, 2016; Blank et al.,, 2003; Hock and Joseph,, 2019), we group them in poles to better compare, categorize, and generalize studies exploring these types of change. Nevertheless, we have included only approaches that consider changes between corpora as this typology was designed for diachronic studies, not changes related to a specific domain or to a specific context (Koch,, 2016; Blank et al.,, 2003; Hock and Joseph,, 2019). In Figure 1 we illustrate the hierachical taxonomy for LSC.
Figure 1: Taxonomy for the poles of Lexical Semantic Change, based on the work of Koch, (2016); Blank et al., (2003); Hock and Joseph, (2019).
We assign studies that track sense broadening or narrowing in the dimension ($\mathcal{D}$) pole and we assume that words that are getting more or fewer senses are increasing or decreasing their dimension value (see Section 5.2). This pole is self-complementary as the change in the value of the dimension addresses both senses. A change in dimension could work as a level of carefulness with translations for example, that broader words should be translated with more care (Mohammed,, 2009).
We regroup in the relation ($\mathcal{R}$) pole changes increasing the relation between one word sense and other senses for this word (see Section 5.3). This kind of changes are commonly related to metaphorization or metonymization, where the meaning of a word depends on the link built by a conceptual (abstract) or material (direct) similarity between concepts. Relations like hypernymy, homonymy, and synonymy are also factors for semantic change (Koch,, 2016; Hock and Joseph,, 2019) as they involve a material relation (related by some cultural association) and can potentially be encompassed in this pole.
For the orientation ($\mathcal{O}$) pole, we group studies that identify the process of amelioration or pejoration or any perceptual change in the ”sentiment” related to words (e.g., criticism, compliment, etc.). This pole also encompasses many forms of change in perception of opinions, as some studies suggest that sentiments like ‘taboo’ (Koch,, 2016) also affect the change in word usage.
To illustrate an example of a word that suffered a change in orientation, and how is it natural to describe it in a single pole, we select the word ‘awful’. In 1850s, it was used with the meaning “full of awe”, and nowadays it means “extremely disagreeable”. As the first sense is semantically closer to good and the second closer to bad sentiment, this word suffered a pejoration as illustrated in Figure 2. Notice that the evolution of the orientation of this word pass first from an amelioration before changing to a pejoration. However, in the same figure, the word ‘awesome’, that has a close relation with ”awe” gained a more positive sense, suffering an amelioration over time.
Figure 2: Change in meaning and orientation for the word awful. In the left side, we reproduce the figure from Hamilton et al., (2016) that shows the evolution of the word ‘awful’ in the embedding space. In the right side, we present the hypothetical function ($\mathfrak{f}$) for this word over time.
In the next section, we will describe our method for reviewing the literature based on the poles presented above.
2.3 Review Criteria
Querying for documents in different databases of scientific papers required an interactive work to include terms that are relevant to our survey. There is no consensus on the terminology used to categorize the papers. Thus, we decided to use a semantic search approach to identify critical papers in the field and enrich our searching query. In Table 1, we present the final set of search terms used to collect articles related to the characterization problem. We used Semantic Search, Scopus and Google Scholar as source of papers. The interactive process has as “end criteria” when the search for the terms does not find any new match. We execute this process for each term identified. The semantic search help us to extend the set of terms. For that, we used an auxiliary software, Research Rabbit [^2] https://researchrabbitapp.com/, to discover similar relevant work.
Table 1: Search terms used to discover articles related to change characterization.
Characterization Search terms
Dimension
“semantic change broadening”,“semantic change narrowing”,
“novel sense change”,“innovative meaning”,“new sense identification”,
“neologism identification”
Relation
“metonymy semantic change”,“metonymization semantic change”,
“metaphor semantic change”,“metaphorization semantic change”
Orientation
“semantic change polarity”,“semantic change feeling”,
“semantic change orientation”,“semantic change emotions”,
“semantic change sentiment”
We gave as input to the software a set of related papers (found in Google Scholar and Semantic Scholar) and we performed an iterative process of adding new papers until reaching a stable set of documents without further improvements. In Figure 3, we present the graph of documents, where green nodes represent the input papers, and blue nodes are papers suggested by the tool after the first interaction.
Figure 3: Graph of selected works (green) and related articles (blue).
We manually filter the papers that characterize word changes based on context, e.g., as a figure of language, but not a fundamental change in the word sense. For example, those approaches that detect if a word was used metaphorically in a sentence, but do not infer if the word suffered a metaphorization over time/corpus.
For relation and orientation, we had a vast range of false positives as many works are trying to capture metaphor or orientation without comparing corpora, so this approach could not infer changes in the sense and were not included in the selected papers.
In the next sections, we investigate in depth three elements of semantic change characterization, (i) the corpora ($T$) analyzed, (ii) the representation method ($\mathfrak{R}$) for capturing characteristics of changes, and (iii) the kind of characterization ($\{\mathcal{D},\mathcal{R},\mathcal{O}\}$) addressed by these methods.
3 Corpora for Semantic Change Analysis
Corpora for semantic change could be selected based on differences in time (diachronic studies), domain (academic fields (Schlechtweg et al.,, 2019)), culture (social media (Carpuat et al.,, 2013; Shoemark et al.,, 2019), countries), language (Takamura et al.,, 2017; Uban et al.,, 2021), and so on, as soon as the pair of corpora ($t_{1},t_{2}$) taken at different contexts have in common a sufficient amount of symbols for comparison.
In diachronic studies, some studies set the problem as two different data distributions with a meaningful gap between the time interval. Some methods divide the corpora into discrete bins or time windows, striking a balance between the size of the window and the quantity of data it contains. In contrast, others create no bin and proceed with a continuous representation of change where the timestamp associated with the text is directly provided to the model (Hu et al.,, 2019; Rosenfeld and Erk,, 2018; Wang et al.,, 2021). These corpora were designed for specific purposes. For example, the first two types of corpus rely on a discretization of word meanings in a comparative perspective, while the third one allows for an evolutionary view of the changes.
With respect to characterization, the collected papers show an increasing interest on diachronic corpora and many of them propose handcrafted rules for automatically annotating the categories of changes (Tang et al.,, 2013). Few works rely on human annotation (Schlechtweg et al.,, 2017). While these approaches have advantages and drawbacks, from a computational perspective, the required data to train representation models is far from being large enough. Recent state-of-the-art models for NLP require training on billions to trillions of tokens (He et al.,, 2021), but diachronic corpora generally present tokens in a scale of millions (Schlechtweg et al.,, 2020). Given that increasing the amount of training corpora is in most cases infeasible, the viable approach to learning semantics in this context is with better algorithms.
Even with the progress in the fields of low-resource learning like rapid adaptation (Neubig and Hu,, 2018), few-shot (Liu et al.,, 2022) or in-context learning (Min et al.,, 2022). There is still a meaningful lack on freely available corpora for semantic change studies (Kutuzov et al.,, 2018).
4 Representation methods for Semantic Change
While interest in detecting semantic changes is growing, there is less efforts being applied on characterizing these changes. Computational linguistics has not yet fully explored the usefulness of having detailed information about a type of change. For example, a change in relation would have a low impact on a sentiment analysis algorithm, while a change in orientation is more likely to affect it.
In this context, the methods to represent meanings are of high relevance to understand what kind of change could be described and characterized. This section investigates in the literature representation methods ($\mathfrak{R}$) based on: word frequency, topics, graphs, and embeddings. We analyzed how each method was used to characterize LSC into the three proposed poles.
4.1 Word frequency
Methods based on word frequency were among the first computational attempts to tackle semantic change and are still at the heart of LSC and of many other linguistic approaches. Analysing word frequency in a corpus contributes to a direct interpretation of change, which in the context of characterization, is generally executed by counting co-occurrence with other words.
4.1.1 Dimension
To detect changes along the dimension axis ($\mathcal{D}$), Sagi et al., (2009) propose a representation method ($\mathfrak{R}$) to count together words and phonemes (e.g. ‘dog’ and ‘do’) and measure the cluster density of the cosine distance on a decomposition of the term frequency matrix. They assume that Latent Semantic Analysis is appropriate to examine vectors representing the context of each occurrence of a given term in a corpus and estimate the cohesiveness by measuring the density with which these token context vectors are related to each other in a vector space. While they explore the methodology for taxonomic changes analysis, they state that the method in fact measures polysemy of words. Tang et al., (2013) propose a framework for semantic change computation by assuming that semantic change is by nature a successive and diachronic development. They further distinguish between word level (i.e. the semantic of a given word) and sense level (i.e. individual sense of a word) and implement their approach based on these two notions. One essential step in their approach aims at computing the senses of a word in a given corpus and then the diachronic distribution of these senses in the corpus. Carpuat et al., (2013) focus on neologism detection in corpora and they apply classification techniques for that. Those techniques are denoted as i) Type-level Features and encompass word frequencies change, the appearance of the word in a new context, and word’s distribution over topics change, ii) Token-level Features that refers to the same kind of features but generalize at the token level. Emms and Jayapal, (2016) propose a generative conditional model to detect if a word gained a new sense over time. This is done by modeling the word sense distribution with Gibbs sampling (a kind of a Markov chain Monte Carlo algorithm for obtaining a sequence of observations). Bochkarev et al., (2022) use a neural network to predict if a word is used as a named entity. By this method, they create a temporal view of usage of a word and they can compare the occurrences of a word to detect if it gained a new usage in the corpora.
4.1.2 Relation
The work of Tang et al., (2013) transforms the problem of relation identification into change pattern detection in time series data. They analyse the semantic change at the word level and the individual-sense level inside a corpora based on rules in order to infer if a word is used as metonymy or metaphor. A different approach computes the entropy of target words (i.e., the unpredictability of the use of these words) to identify if a word is being used in a non-regular context Schlechtweg et al., (2017). The higher the entropy of a word is, the more likely this word is used in a newly created metaphoric sense. They are using a corpus of German document to validate their approach.
4.1.3 Orientation
Frequency metrics to detect the co-occurrences with sets of ”good-” or ”bad-” sentiment words has been widely used to infer the amelioration and pejoration of a word over time, especially for neologism (Cook and Stevenson,, 2010; Jatowt and Duh,, 2014). Similarly, Buechel et al., (2016) use seed words in German and English to track the emotions captured in corpora. Additionally, the authors represent it in the Valence-Arousal-Dominance framework instead of following a negative and positive connotation. A limitation of the later study is that it does not track the evolution of orientation in words.
Frequency approaches are, for a long time known to be limited for the analysis of semantic of words, given their inability to capture similarity, as stated in the work of (Mikolov et al.,, 2013), and compositionality, as pointed out by (Mitchell and Lapata,, 2010). For instance, documents using different words could describe the same thing. The n-gram approach, to some extent (Cheng et al.,, 2006), improves compositionality understanding but is hindered by the statistical significance or Zipf distribution (Zipf,, 1949), which is outperformed by complete sentence understanding as showed by Socher et al., 2013a .
Another limitation of frequency-based methods is its inability to represent sense and a change in word usage could not be differentiated between sense broadening or narrowing. Therefore, an extra step is required to deal with polysemy. Some authors tried normalizing frequency (Hamilton et al.,, 2016), however, as normalization developed for one corpus is hardly applicable to another corpus (Brezina,, 2018), this method fell into disuse (Kutuzov et al.,, 2018). Another approach that fell into disuse is the definition of general laws for semantic change (Dubossarsky et al.,, 2017). The use of co-occurrence tables for a the analysis of semantic change misjudged polysemy and prototypicality. Deep study of co-occurrence tables concluded that there is no statistical correlation between frequency with polysemy or semantic change.
Finally, frequency-based methods has shown low performance to detect orientation of words (Socher et al., 2013b, ) as well as metaphor and metonymy. Since the words are deeply related to the used context, the lack of compositionality could mismatch irony or negation, and could not recognize the right ”relation” between word meanings.
4.2 Topics
Topic modeling is a statistical method used to identify the most important themes addressed in a collection of documents. This is done by grouping words that are frequently used together under topics. For example, a topic model might identify the topic of “bank” in a corpus of news articles by grouping together words like “government,” “money,” and “insurance”.
Semantic analysis is a type of intelligent algorithm that uses topic modeling to understand the meaning of words and phrases. It does this by tracking the different topics that a word is associated with in a context including time, domain, culture, etc. For example, a topic model could be used to understand the different meanings of the word “bank” in the context of finance (i.e., a financial establishment) and in the context of a river (i.e., the land alongside a body of water), or the word “compute” from its early days as a large mainframe machine to its current meaning as a small and personal device. The different topics associated with a word are tracked and compared in order to identify changes in their senses and, when possible, characterize these changes.
4.2.1 Dimension
These approaches assume that changes in the number of topics assigned to a word can indicate changes in dimension. Lau et al., (2012) introduced the idea of novelty as the proportion of usages of a given word corresponding to senses in the focus corpus and reference corpus. The goal is to use word sense induction based on topic modeling to detect if a word is used with a new meaning. Cook et al., (2013) extended the work of (Lau et al.,, 2012) by quantifying the use of a topic and estimating that a “cultural silent” topic that becomes frequent in the focus corpus has high probability to have neologisms. For that, they introduced the concept of relevance of a topic and compute a rank of potential neologisms. Lau et al., (2014) improved the definition of novelty by adding the frequency of the word in the corpus, then they proposed to combine novelty and relevance to rank the neologisms.
4.2.2 Relation
To the best of our knowledge, no methodology for tackling conceptual change with topic-based representation has been published, while some works use it in a synchronic detection setting (Heintz et al.,, 2013).
4.2.3 Orientation
To the best of our knowledge, no methodology approaching orientation change with topic-based representation was published. Some works in sentiment analysis literature use topic models in a workflow in order to refine the data used by a classifier to predict the perceived sentiment. However, topic models are not directly applied for the classification task (Wankhade et al.,, 2022).
Topic modeling methods are powerful tools for text analysis, but it is important to be aware of the challenges associated with interpreting the results. The outcomes often require manual analysis to consolidate and to give sense to the topics. This is because the aggregation of words are statistically defined and can be difficult to be understood by humans, which can lead to misinterpretations of word senses.
By using a variety of evaluation metrics (such as topic coherence and topic diversity) and visualization techniques, the manual work can be reduced. In the cited approaches, the authors automatize the selection of topics by performing posterior filtering based on the similarity with dictionary definitions. Although the potential of this method to detect neologisms, its use for change characterization still need deepen investigations.
4.3 Graphs
Graph structures are a natural way to represent relationships between words. They are often used in semantic change research, where they are typically constructed manually based on dictionaries and human supervision. However, graph structures can also be constructed automatically based on word co-occurrence. Co-occurrence graphs are better at capturing context than count tables because their edges represent more complex relationships, such as cliques of words that share some context. For example, the words ”feline”, ”cat”, and ”pet” would be likely to form a clique in a co-occurrence graph, indicating that these words are often used together and have close meanings.
4.3.1 Dimension
Comparing the size of a graph over time can provide rich information about the semantic change of a word. Biemann, (2006) was the first to implement this task in order to detect the sense broadening of words. They created a graph of associated nouns that occur in corpora, then applied the Chinese whispers algorithm to find clusters in the graph. By comparing these clusters over time, they were able to infer whether a word had acquired a new meaning. Mitra et al., (2015) analysed the Tweets and count the word co-occurrence based on bi-grams as a way to identify potential relationships. The graph was built based on these relationships and they run a graph-clustering algorithm and Chinese Whispers to determine the set of senses. Each cluster is considered as a different sense of a word. By comparing this sets of senses over time they can infer if a word split its senses or the senses becomes broader (joined by other words). Mathew et al., (2017) improved filtering and sense selection over previous methods introduced by (Mitra et al.,, 2015; McCarthy et al.,, 2004; Lau et al.,, 2014), to better capture sense evolution in Google Books and newspaper datasets. Tahmasebi and Risse, (2017) consider the graph of occurrences of noun phrases with length two words as a cluster of meaning, then they merge sub-graphs based on the lin similarity. They track dimension by analysing how this sub-graphs change over time. Ehmüller et al., (2020) build a co-occurrence graph, filtering with point-wise mutual information to remove non relevant edges, and then create an ego-network to later run graph-clustering algorithms. This approach allowed to build a sense tree333a tree of word usage evolution, where new senses are new branches. Similar to previous work, the authors restricted their analysis to nouns only.
4.3.2 Relation
Graphs are a good way to represent word relations. Some approaches have used graphs to detect metonymy usage Teraoka, (2016); Pedinotti and Lenci, (2020). Their model uses the graph-based distributional model of event knowledge to determine a contextualized representation of word semantics, which is obtained by integrating the lexical embedding of a word with a vector representation of the expectations activated by the context for the event role of the word. However, to the best of our knowledge, no work proposed to use graphs in a representation level to compare which words have changed their relations between corpora.
4.3.3 Orientation
The orientation of words generally is inferred from similarity or proximity with respect to seed words. Many word graphs present this construction and some works (Reagan et al.,, 2017; Barnes et al.,, 2021) use these graphs to infer orientation. However, to the best of our knowledge, no work used graphs to compare the orientation of words in different contexts.
Current graph-based approaches did not explore modern constructions of graph structure like knowledge graphs or graph embeddings to detect and characterise LSC. This kind of representation could help inferring changes in the proposed poles, e.g., Wikidata has information about hierarchical relations between entities, e.g., Washington capitalOf U.S.A., that can be useful to construct word material relation and consequently the analysis of metonymization.
Wordnet, Babelnet and Conceptnet are examples of graph structures used auxiliary to many methodologies for change detection (Giulianelli et al.,, 2023; Tang et al.,, 2023). However, these structures are under-explored (generally used only as a sense storage), as they lack temporal information, like timestamps for first occurrence.
4.4 Embeddings
Word embeddings are representations of words as numeric vector. The vector is built in such a way that words that have close values in the vector space are expected to have similar meanings. The distance between these words can help the interpretation of their sense.
Word embeddings have been used for a variety of tasks, including LSC identification and characterization. This is done by comparing word vectors between two moment in time, the distance between them and with respect to the neighboring vectors can indicate if the sense of the word changed (Hu et al.,, 2019).
4.4.1 Dimension
Ryskina et al., (2020) propose a statistical approach to identify neologism based on frequency count and semantic sparsity (measured with embeddings). They claim that words and concepts suffer from supply and demand curves, so emerging concepts (i.e., neologism or sense broadening) would appear to populate the sparse semantic region until the space becomes uniformly covered. In the work of Vijayarani and Geetha, (2020), the authors demonstrated that vector embeddings and lexical features can be used to detect sense narrowing. They explain with the characterization of the changes in meaning of the word “tablet”, in English. They found that the word had undergone a narrowing of meaning because now it was commonly used to describe a portable computer than rectangular shaped objects (as it was its initial utilisation). By analysing the deviation of a Gaussian word embedding, Yüksel et al., (2021) identified a way to measure the narrowing and broadening in word senses. In Moss, (2020), the authors also performed a training per sense of these embeddings.
4.4.2 Relation
Computational analysis was used by Fonteyn and Manjavacas, (2021) to investigate the process of metaphorization of the concept “to death”. They combine methods of cluster analysis and sentiment analysis to contextualize the occurrences of this phrasal verb and detect if a metaphorization has taken place. Another approach is to create a training corpus and use it to train a BERT model to predict if a word is a metaphor or literal. The work of Maudslay and Teufel, (2022) follows this approach and the authors propose a framework to detect words that started being used as a metaphor.
4.4.3 Orientation
Methods based on word embeddings have been widely used for many sentiment analysis tasks (Raffel et al.,, 2019) and to investigate sentiment at lexicon level (Ito et al.,, 2020). In the context of LSC characterization, Fonteyn and Manjavacas, (2021) uses the distance between the word vectors of ‘good’ and ‘bad’ as a means to measure polarity in the expression ‘to death’.
In general, word embeddings have an unstable and many times incoherent behavior, as pointed out by (Hellrich and Hahn,, 2016; Gladkova and Drozd,, 2016; Günther et al.,, 2019). The reason is that vectors representing words with the same meaning could be sensitive to minor variations in the corpora or in the architecture utilized for language modeling, thus different vectors can represent the same meaning in different corpora (Wendlandt et al.,, 2018; Leszczynski et al.,, 2020; Mickus et al.,, 2020). This undesirable effect is often observed for low frequency words in the corpus, which corresponds to most words in a vocabulary (Zipf,, 1949), and particularly for those with polysemy.
To mitigate the stability problem, where different training routing led to very different learned vectors, authors recurred to larger embedding sizes, quantization (May et al.,, 2019), or alignment between trained embeddings (Hamilton et al.,, 2016). However, the vector representation is already a burden, as standard approaches to semantic change identification suffer from meaning conflation deficiency (Camacho-Collados and Pilehvar,, 2018).
The conflation deficiency happens when a polysemous word is represented by a single vector. Suppose that a word $w$ in moment $t_{1}$ has two sense, e.g., $w_{t_{1}}=\{s_{1},s_{2}\}$, where the dominant sense (the most common sense in the corpora we are looking at) is $s_{1}$ and this configuration pushes the embedding vector to its semantic region. At a later moment in time $t_{2}$, the word still have both senses, e.g., $w_{t_{2}}=\{s_{1},s_{2}\}$, but the dominant sense is now $s_{2}$. The result is that the embedding vector will shift towards the semantic region of the concept $s_{2}$. There were no real semantic changes in the word. But, the example can misled semantic change detection algorithms to this wrong conclusion.
Some works tried to address conflation deficiency by adding a vector for each sense (Neelakantan et al.,, 2014; Liu et al.,, 2015), which generally requires additional annotation or steps to discriminate senses (i.e., given a word, which vector will be used).
An extension of these works includes contextualized words, and create embeddings based on pre-trained language models, generally extracted from ELMO (Peters et al.,, 2018) or BERT (Devlin et al.,, 2019). The main difference is that instead of using the pre-trained embeddings for each word, this approach uses the outputs of the transformer’s final (contextualized) layers to capture meaning. These outputs are then clustered or compared using cosine similarity to measure the sense of the word. For instance, the representation of the word ‘mouse’ in a text will be assigned to the correct semantic region, such as ‘animal’ or ‘computer device’.
Nevertheless, while they presented an improvement to sense detection, they produce words representation that are too contextual. Thus, words with the same meaning can have very different vectors if they are in a different context, and the representations are anisotropic. The distribution in the semantic space is not uniform, leading to non-interpretable distances in this space (Ethayarajh,, 2019). So for diachronic studies, this methodology is not reliable as it gives many false positives (Kutuzov et al.,, 2022).
Recent studies try to improve the isotropic orientation and sense similarity (instead of contextual similarity) through contrastive learning, using entailment (Gao et al.,, 2021; Reimers and Gurevych,, 2019) or pre-training on similarity tasks (Zhou et al.,, 2023). With a pre-training objective of reducing the distance between similar senses in the semantic space, these works improved pre-trained textual and sentence similarity models.
Another set of approaches attempts to model the polysemy of words directly. In the characterization literature, some works used density-based embeddings (word2gauss) to track this behavior. In Moss, (2020), the author selected some properties of this embedding to measure the dimension of change. A more general review of embeddings capabilities and their limitations is presented in Apidianaki, (2022).
Current state-of-the-art methods in LSC Identification rely on fine-tuning cross-language models trained on Word Sense Disambiguation (WSD) tasks for English data. These models are then used to extract word embeddings. This approach leverages the phenomenon of transferability between languages, where English data can improve the model’s performance on tasks in other languages, such as Russian (Pivovarova and Kutuzov,, 2021) and Spanish (Zamora-Reina et al.,, 2022) tasks.
4.5 Summary
Table 2: Comparison table between studies highlighting the type of methods utilized and category of change. We mark which kind of characterization the work conducts $\mathcal{D}$ for dimension, $\mathcal{R}$ for relation, and $\mathcal{O}$ for orientation. Also, we highlight the main representation method ($\mathfrak{R}$) for the word meaning.
[…]
Biemann, (2006) X - - Graph
Sagi et al., (2009) X - - Frequency
Lau et al., (2012) X - - Topics
Cook et al., (2013) X - - Topics
Tang et al., (2013) X X - Frequency
Carpuat et al., (2013) X - - Frequency
Lau et al., (2014) X - - Topics
Mitra et al., (2015) X - - Graph
Emms and Jayapal, (2016) X - - Frequency
Mathew et al., (2017) X - - Graph and Topics
Tahmasebi and Risse, (2017) X - - Graph
Ryskina et al., (2020) X - - Embeddings
Vijayarani and Geetha, (2020) X - - Embeddings
Moss, (2020) X - - Embeddings
Ehmüller et al., (2020) X - - Graph
Yüksel et al., (2021) X - - Embeddings
Bochkarev et al., (2022) X - - Frequency
Schlechtweg et al., (2017) - X - Frequency
Fonteyn and Manjavacas, (2021) - X X Embeddings
Maudslay and Teufel, (2022) - X - Embeddings
Cook and Stevenson, (2010) - - X Frequency
Jatowt and Duh, (2014) - - X Frequency
Buechel et al., (2016) - - X Frequency
Table 2 summarizes the literature survey conducted in the previous sections, for semantic change characterization. In the rows we add the most relevant papers, while the columns are the main properties of LSC evaluated in this survey. The reference to the papers is in the first column. The second, third, and forth columns represent the three poles of change characterization, i.e. dimension ($\mathcal{D}$), relation ($\mathcal{R}$), orientation ($\mathcal{O}$). Finally, the technology adopted to compute and/or represent the sense ($\mathfrak{R}$) is in the fifth column. Notice that an “X” in a cell indicates that the paper applied the sense computation technology (listed in the fifth column) to analyze the corresponding “pole”.
In the table, we can observe that no methodology was used to characterize all three poles, showing the complexity of this problem and the preference of the authors for some specific problem. The majority of the studies were developed to characterize changes in the dimension pole. It seems to be the pole where the computational linguistic approaches get the best results. Another observation from the table that deserve our attention is the still under-explored use of graphs and topics methods to represent sense, compared to embeddings and frequency. It is also important to notice that the most recent works use embeddings, showing the current trends to this kind of approach.
Although computational linguistics has been used to characterize Lexical Semantic Change (LSC) for over ten years (the first published work was in 2006, according to this survey), there is still no standardized formalization of these changes. Each author proposes their own solution, and many of the proposed interpretations of change are presented in the papers in a textual or semi-formal format. To contribute to the discussions and to the definition of semantic change characterization, we propose a set of formal descriptions of change characterization and provide concrete examples of the subtasks required to cope with the LSC problem in the next section.
5 Formalization
In computer science, formalizing a problem is an essential step. It allows us to reason about the problem using computational concepts and directly design algorithms to solve it. This process is particularly crucial in research fields. A well-defined formalization ensures that researchers are working with the same underlying assumptions and terminology. This shared ground facilitates the development and comparison of different methodologies. In our review, we observed that while the analyzed papers conceptually addressed the same problem, the lack of a common formalization made it difficult to compare them effectively. To address this gap, this section proposes a formalization based on the reviewed studies and insights gleaned from linguistic literature to guide future research in this area.
However, it’s important to acknowledge the limitations of purely computational approaches. Linguistic literature, exemplified by Campbell, (2013, p. 221-246), explores semantic change through historical and socio-cultural lenses, examining how societal events influence word evolution. Replicating this approach entirely through computational analysis remains an open challenge. Therefore, we focus on formalizing methods based solely on semasiological comparative analysis, which compares the meaning of words across time without relying on additional knowledge of the cultural context in which those meanings were produced.
In statistics (and machine learning), the problem of drift distribution is formally defined (Gama et al.,, 2014). However, in linguistics, the definition of word’s meaning (from a “semasiological” perspective) still relies on ideas such as ‘concepts’ and ‘sentiments’ without providing objective definitions for this terms (Koch,, 2016; Traugott,, 2017; Blank et al.,, 2003), making a formal definition of drift distribution more challenging. This vague notion of change complicates the implementation of procedural methodologies to analyse, compare, and evaluate LSC.
We build our formalization by posing the meaning of a word as as a set of glosses, inspired by Hock and Joseph, (2019, p. 194), and define all constructions in set theory. First, we assume that for a set of corpora of text $T$ there is a universe of senses $S_{T}$, that holds all the senses for that corpus. Given an alphabet $\alpha$, we define $\mathcal{S}$ as a function that takes a string of the closure $\alpha$ where $\alpha^{*}=V$ (the vocabulary) and returns a subset of $S_{T}$. For instance, given a word $w\in V$ and a corpus $t\in T$ , the subset of senses $S(w,t)$ for the universe of senses ($S_{t}$) of the word $w$ in this corpus is the following:
$\begin{split}\mathcal{S}:V\times T\rightarrow\wp(S_{t}),\\ \mathcal{S}(w,t)=\{s_{1},s_{2},...,s_{k}\}.\end{split}$ (1)
We define that two sets of senses are different if, and only if, the symmetrical difference between a set of senses for the word $a$ in a corpora $x$ and the word $b$ in a corpora $y$ is non-empty, where $a,b\in V$ and $x,y\in T$
$\begin{split}\mathcal{S}(a,x)&\neq\mathcal{S}(b,y)\iff\\ (\mathcal{S}(a,x)-\mathcal{S}(b,y)\neq\emptyset)&\lor(\mathcal{S}(b,y)-\mathcal{S}(a,x)\neq\emptyset),\end{split}$ (2)
or in simple terms, they are different if the sets of senses from corpus $t_{1}$ and $t_{2}$ have different elements.
5.1 Semantic Change
Koch, (2016, p. 23) presents a definition of meaning change that combines semasiological and onomasiological perspective. In Figure 4, ‘belly’ was initially associated with the source concept ‘bag’ (M1), later it acquired the meaning of the target concept ‘body part between the breast and the thighs’ (M2). However, in the field of computational linguistics, we primarily focus on the semasiological view. Here, our interest lies solely in the word itself and the evolution of its meaning, rather than delving into the origins of the concepts it represents.
Figure 4: Definition of semantic change adapted from Koch, (2016, p. 23,25).
The process of meaning change is depicted on the right side of the figure. At time T1, the word ”belly” possessed only the meaning M1. Subsequently, at T2, it began to be used with the new meaning M2 as well. This usage of M2 increased over time (T3-T5) until it eventually became the dominant meaning (T6). Finally, at T7, the original meaning M1 was completely lost.
To formalize this definition, we say that a word $w$ had a semantic change ($\mathcal{C}$), if the set of senses, given by the function $\mathcal{S}$, evaluated in the corpus $t_{1}$ is different from the set of senses in $t_{2}$, given that $t_{1}\neq t_{2}$,
$\mathcal{C}(w,t_{1},t_{2})=\begin{cases}\text{True}&\mathcal{S}(w,t_{1})\neq\mathcal{S}(w,t_{2})\\ \text{False}&\text{otherwise.}\end{cases}$ (3)
The subset of words in vocabulary $\bar{W}\subset V$ that suffered a semantic change between corpus $t_{1}$ and $t_{2}$ is given by,
$\bar{W}_{t_{1},t_{2}}=\{w\mid\mathcal{C}(w,t_{1},t_{2}),\hskip 10.00002pt\forall w\in V\}.$ (4)
These changes in lexical meaning between corpus are caused by many different phenomena, like time period (1930 and 2020, gay), culture (Brazil and Portugal, rapariga), and domain (engineering and biology, cell).
5.2 Change in Dimension
We define that a word $w$ had a changed in dimension if the set of meanings increased or decreased, with respect to the cardinality of the set, when comparing two corpus $t_{1},t_{2}$. It means that the word suffered a broadening or a narrowing change between these corpora.
$|\mathcal{S}(w,t_{1})|\neq|\mathcal{S}(w,t_{2})|,\hskip 10.00002ptw\in V.$ (5)
This definition is compatible with Koch, (2016, p. 24-26) concept of innovative/reductive meaning change, in which words acquire new senses (sememes) by being used with different meaning. However, it differs for the more general broadening notion, where words that undergo broadening may gain extension, meaning that they are used in a broader sense than initially, but they may also lose intention, which is the precision of the word’s meaning, making it more vague.
Under this setting, split, birth and join with respect to a single lexical are particular cases of change in dimension, e.g., birth or death is considered a change from or to an empty set. Split and join are also instances of broadening or narrowing with respect to a word. For instance, the word ‘plane’ after the 19th century gained the sense of an aeroplane, while previously was only used to describe the act of planing, or a flat surface. This new sense greatly increased the dimension of the word plane, as many correlated senses, like “travel in an aeroplane”, was also added to plane senses.
In this survey, we define broadening as the degree of polysemy in a word. This definition is aligned with the NLP literature, given its application in data analysis and its measurable indicator of validity, compared to the notion of intention. We differ from the notion defined in Tang et al., (2013), where broadening is a superclass of metaphorization and metonymization, as the literature reviewed here does not perform this aggregation.
5.3 Change in Relation
The phenomenon of metaphorization or metonymization occurs when a word takes on a new meaning that, to some extent, inherits qualities from its original meaning through a figurative relationship the speaker aims to convey (Lakoff and Johnson,, 2008, p. 3-6). The original meaning is invoked by establishing a material or abstract relation with the initial sense, where the listener is able to solve the non-literal target concept. In the realm of human cognition, material relations and abstract relations are distinguished by the type of association formed between concepts. Material relations are established through the creation of contiguity associations in the physical world. This means that the relationship between the target and source concepts is built upon their physical proximity or interaction. In contrast, abstract relations are formed through similarity associations within the conceptual world. Here, the connection between concepts arises from shared characteristics or underlying principles, independent of any physical manifestation.
A material relation is built when we recur to some cultural association established between the concepts (source and target) that are interleaved under a shared context (a frame) (Koch,, 2016, p. 42-51), e.g., in the sentence “The proposition made by Washington was rejected by the other countries”, we are replacing the location of the institution by the institution itself. Other forms of metonymy use cause and effect, contiguity, instrumentality, possession, etc [^4] Littlemore and Tagg, (2018) group the forms of metonymy as ‘referential,’ ‘propositional,’ ‘illocutionary,’ and ‘situational.’.
On the other hand, an abstract relation is achieved by associating properties that evoke sensory experiences, such as emotions, shapes, dimensions, positions, and more (Lakoff and Johnson,, 2008, p. 56-60). For example, “The arm of the chair is broken,” in this sentence the listener imagine the chair as a human being and associate body parts to relative positions of the chair. In this case, the literal sense is elucidated by examining the figurative properties that the listener can grasp within the sense itself.
The theory on how metonymies and metaphors are defined is informal, where generally relying on subjective perception, such as personal experience, or definitions that overlap with other figurative modes. For example, Choi et al., (2021) defines metaphor as “literal meaning of a word is different from its contextual meaning”, which is also the case for metonymy. Many authors in linguistic try to disambiguate the relation relying on imprecise definitions (Campbell,, 2013, p.224-225), and as discussed by Koch, (2016, p. 42-51) there’s no precise rule to define a metonymy. In Schlechtweg et al., (2017) the authors only measure the entropy of a word as a means to capture metaphorization, it does not provide evidences that this measure does not vary for the case of metonymization or a new sense not related to the previous senses  [^5] For example, the word ‘apple’ used in a technology context increases entropy but is not a metaphorization process, but a new sense..
Given the above considerations, we define that a word $w$ had changed in relation, if the aggregated sum of the relational similarity ($\mathit{l}:S\times S\rightarrow\mathbf{R}$) between senses in $\mathcal{R}(w,t_{1})$ changed for $\mathcal{R}(w,t_{2})$. Where $\mathit{l}$ takes two senses $s_{i},s_{j}$ and returns a real value, that correspond to the material or abstract relation between these senses (greater the value, more related the senses are), i.e.,
$\mathcal{R}(w,t)=\left\{\begin{aligned} \sum^{N-1}_{i=1}\sum^{N}_{j=i+1}{\mathit{l}(s_{i},s_{j})}\hskip 30.00005pts_{i,j}\in\mathcal{S}(w,t)\\ 0\hskip 60.00009ptN\leq 1,\end{aligned}\right.$ (6)
where $N=|\mathcal{S}(w,t)|$.
Given a word $w$, we define that it increased in relation with respect to corpus $t_{1},t_{2}$, if $\mathcal{R}(w,t_{1})<\mathcal{R}(w,t_{2})$, and decreased if the opposite happens.
While we have no methods to perfectly differentiate the material and abstract relation, there’s no excluding factor (a word can suffer metaphorization or metonymization at the same time), and for many applications the relevance is tracking if a word is being used with the literal or figurative meaning. We decided to simplify the interpretation of these relations as following: whether it increases in material or abstract relation, we consider it as ”increasing in relation”, if a new sense a word acquires is related to the original sense, and “decreased in relation”, if it is less used in a figurative manner.
5.4 Change in Orientation
The change in word’s orientation (positive or negative connotation) reflects the amelioration or pejoration in the usage of a word from one corpus to another. This is measured by analyzing how the collective perception of a word’s meaning evolves within a specific context (Campbell,, 2013, p. 227-229). A classical example is the word ‘terrific’. Originally, ‘terrific’ meant ‘terrible’, carrying a negative connotation. Then, its meaning shifted to ‘formidable’, signifying something positive and impressive.
Buechel et al., (2016) uses a Valence-Arousal-Dominance setting, where the valence axis represent positive and negative sentiment, but it also extends the emotion perception to more dimensions, arousal and dominance. While this in-depth analysis has its value, it is not explored in the linguistic literature and the majority of papers in sentiment analysis (Cui et al.,, 2023), so we restrict our formalism to the more general polarity setting(positive/negative), to be aligned with the current research and applications (Susanto et al.,, 2020).
We define that a word $w$ had changed in orientation, if for some notion of orientation $\mathfrak{f}$ (feeling), there is a mapping from a word $w$ in a corpora $t$ to an orientation value, for example, in case of positive, neutral, or negative feeling,
$\mathfrak{f}:V\times T\rightarrow\{-1,0,+1\},$ (7)
the orientation changes between corpora $t_{1},t_{2}$,
$\mathfrak{f}(w,t_{1})\neq\mathfrak{f}(w,t_{2}).$ (8)
Analogous to the previous definition, if $\mathfrak{f}(w,t_{1})<\mathfrak{f}(w,t_{2})$ it suffered an amelioration, and a pejoration if $\mathfrak{f}(w,t_{1})>\mathfrak{f}(w,t_{2})$.
In this task, we rely on a subjective notion of orientation addressed by a function ($\mathfrak{f}$) that is arbitrarily defined. Generally, studies define ‘orientation’ as the proximity to seed words with a positive or negative connotation (a polarity setting) manually selected to keep the same orientation over different corpora. However, the orientation problem is broader than just polarity, as we can classify words in more aspects than positive and negative feelings.
5.5 Semantic Change Identification
Consider a function $f_{\mathcal{C}}$ that takes two representations $\mathfrak{R}$ (e.g., embeddings, graphs, etc.) of the word $w$ in corpora $t_{1},t_{2}$. We define the problem of semantic change identification as assigning a value ($y$) to this word,
$f_{\mathcal{C}}(\mathfrak{R}(w,t_{1}),\mathfrak{R}(w,t_{2}))\rightarrow y.$ (9)
Schlechtweg and im Walde, (2020) proposes two notions of semantic change, the binary classification, e.g., $y\in\{0,1\}$, and the graded (continuous) semantic change, e.g., $y\in\mathbf{R}$, where $y$ is a distance between word representations [^6] While in Schlechtweg and im Walde, (2020) they consider the graded change for a particular purpose, i.e., the Jasen-Shannon distance, we generalize $y$ to any distance measure applicable to the representation of words..
The binary classification task is aligned to the classical LSC literature (Blank,, 2012, p. 113) and to the human perception which in many cases are influenced by a sensory threshold (Smith,, 2008, p. 34). This notion of change is interested on the gain/lost of senses in a word. There’s no general threshold criteria for the binary setting, it is not enough to have different representations, but the magnitude of this difference should be aligned to human judgment. For example, in Schlechtweg et al., (2020) they define a criteria for the graph representation based on corpus size. However, there’s no measure on how it correlates with human perception of change and how it compares to other representations.
The graded task is originated from the computational linguistics literature, as in computers a natural way to compare objects is measuring the distance between its representations. This approach for semantic change allows a more in depth analysis of change as the grade of change can be compared across a wide range of contexts (history, cultures, etc.).
5.6 Semantic Change Characterization
Deriving from the reviewed literature, we define semantic change characterization as a classification problem where we need to assign identified semantic changes to categories, such as $f_{\mathcal{D}},f_{\mathcal{R}},f_{\mathcal{O}}$,
[…]
While the first equation only measured some notion of distance between the representations, for characterization we first apply the type function on the representation, to then compare. For example in Buechel et al., (2016) we have a orientation characterization, e.g.,
$f(\mathcal{O}(\mathfrak{R}(w,t_{1})),\mathcal{O}(\mathfrak{R}(w,t_{2})))\rightarrow y,$ (11)
where the $\mathcal{O}$ function is the Valence-Arosal-Domance framework and they measure the distance $f$ through the $\beta$-coefficient of a linear regression.
Again we set two possibilities of change, binary and graded. In this setting the binary form allow us to answer interesting questions in an automatic manner, for example, in which point in time the word ‘head’ started being used in a figurative manner? Another possibility is to track the precise corpus were words pejorated, and then manually understand the socio-linguistic process to this phenomena, saving many hours of historiography work.
6 Discussion
In this section, we synthesize the reviewed approaches for change characterization and discuss their limitations according to the dimension, relation and orientation aspects previously introduced and formalized.
6.1 Dimension
The majority of the studies found during this survey propose methods and tools to characterize how words change in dimension, while the other poles were less studied. Change in dimension seems to be easier to detect than the other categories because it does not rely on subjective interpretation, such as relation or feeling, but mainly on countable quantity (see Section 5). Works along this axis often rely on word sense induction, a research area with well developed tooling and many data sources, such as dictionaries and synsets, used to implement supervised learning.
A major characteristic that differentiate the approaches is the word representation method. We identified four main types of word representation methods, used to study dimension changes in LSC. These methods have different trade-offs in implementation cost, representation power, and interpretability. It is difficult to compare the effectiveness of each method for this state-of-the-art, given the qualitative and exploratory nature of the reported results, the use of different corpora (not always available), and the need for external knowledge to acquire meaning (difficult to reproduce). However, the analysis reveals that word frequency and embeddings-based methods produce better results to identify neologism (new words), while topics and graph-based approaches are more effective at detecting sememes (new senses) over time. A limitation for the approaches implemented based on dictionaries is that the meaning of new words are not always in the dictionaries because there is a temporal delay between the first occurrences of these neologisms and their publication in dictionaries. Another limitation comes with the presence, in dictionaries, of metaphorical senses at the same level as the literal ones, increasing the complexity to make the distinction between ‘dimension’ and ‘relation’. Some authors prefer to group these two poles into one, as in (Tang et al.,, 2013) . The limitation for topics- and graph-based approaches is that the interpretation of the meaning of a topic (or a branch of a graph) is not always clear and can require the intervention of an expert to, for instance, to regroup topics/branches with the same meaning and/or to map these groups to word’s senses.
6.2 Relation
Outside of this LSC survey, there is a vast body of work proposing methods and tools to predict if a word in a sentence has a metaphorical sense (Heintz et al.,, 2013; Shutova and Sun,, 2013; Hovy et al.,, 2013). These approaches analyze texts, but they do not consider the temporal information, thus they cannot express if a word suffers from metaphorization over time. Some works claim to detect metaphorization of words (e.g., (Tredici et al.,, 2016) ), but they lay on the general hypothesis that changing sense of words implies in a metaphorical change, which still needs to be demonstrated.
For metonymization, we could not find studies that performed metonymy identification in a comparative setting. However, related recent studies investigate the usage of metonymy within different methods to build the material relation identification (Nastase et al.,, 2012; Soares et al.,, 2019; Mathews and Strube,, 2021). We expect that future investigations can extend these approaches to allow detecting metonymization of words.
Among the four meaning representation methods analyzed, frequency-based approaches were the first to be investigated and provided quite good outcomes. However, the hypothesis used by the authors to define a relation is quite similar with the one for dimension, what makes the interpretation of the ‘good outcomes’ more difficult. Embedding is gaining the attention of the community and in recent work, this kind of method has be widely adopted, but the correct interpretation of the published outcomes suffers from the same difficulty observed for the frequency-based approaches. Although the known quality of graphs to well represent relations, this method is still under-investigated to purpose of relation change identification.
(Tang et al.,, 2013) mainly rely on dictionaries of heuristics to infer metaphoricity, which served for qualitative analysis, but still require more extensive validation. A single work used ground-truth annotations for metaphors (Schlechtweg et al.,, 2017), but did not reported a final precision of the proposed methodology.
As noticed, studies about the ‘relation pole’ are few and not conclusive. The low interest of the research community on this topic could be potentially justified by three reasons: (I) The difficulty to formally define, from a computational perspective, what metaphor and metonym are. (II) The rare and insufficient quantity of annotated corpus for relation detection. (III) The nonexistence of standard metrics to measure (and compare) approaches dealing with the detection and characterization of material and abstract relations.
6.3 Orientation
The interest in orientation analysis is increasing with the wide adoption of social networks as media for advertising products and companies. Orientation is usually determined by using a feeling function ($\mathfrak{f}$) that relies on seed words and on the assumption that this seed words do not vary over time. As we illustrate in this survey, the meaning of words evolve over time and can be different between communities. The example of the word sick shows that it can have two opposite orientation according to the context of use. Creating corpus that include (with statistical significance) all these variations is a complex task. In general, we can observe that the annotation process could suffer from many biases, both from the annotators and from the data. The selected seed words could not generalize to be applied to all contexts or in long moments in time. Moreover, a series of misleading results could incur based on seed words, as presented in Antoniak and Mimno, (2021).
From the analysis of the four methods used to represent the meaning, embeddings and frequency-based approaches are the most commonly adopted. Few works use graphs and we could not find any approach that uses topic analysis for detecting the orientation of words. The main reason for the preference for embedding and frequency methods is that current approaches compute the distances or co-occurrence of a word with seed words. It is easier and faster to do this using embedding and frequency-based approaches. Graphs and topics are not designed to determine this and require more complex reasoning and assumptions.
Although there is a huge body of work proposing solutions for analysing sentiment at sentence level, fewer of them try to infer sentiments at a word-level (Guo et al.,, 2021), and in the word-level, only a small fraction is performed on a comparative manner, as we accessed in this survey.
An important outcome of this survey is that there is an harmonization on how change in feeling are defined and applied. However, we could not observe that there is also an harmonization on how the error analysis for the feeling function is implemented. Some works adopt binary orientation function (positive, negative), while other multidimensional orientations (valence, arousal, dominance) (Buechel et al.,, 2016). There is still further study to analyze which function ($\mathfrak{f}$) is more suitable for characterizing change.
One can observe some converging points from the many approaches analyzed in this survey. This is the case, for instance, for the difficulty to evaluate and compare the obtained results because, for most of the tasks, no gold standard exists. Many papers have raised this problem (Schlechtweg et al.,, 2017), but some initiatives to create reference datasets are under development.
Another interesting point is the multilingualism aspect. In the vast majority of cases, only English documents are used for the experiments, which makes it difficult to generalize the proposed methods to other languages with different roots and grammatical rules.
The recent emergence of new-generation of language models has opened up new possibilities for the identification and characterization of semantic changes. These models’ ability to handle large corpora of documents and to deal with contextual information expressed in different languages is important for the analysis of semantic drift. However, this area of research is still under-explored.
7 Conclusion
In this survey we reviewed current works on semantic change characterization, with respect to the mainstream typology, (broadening, narrowing, amelioration, pejoration, metaphorization and metonymization). We observed four major approaches to represent meaning (frequency, topics, embeddings, and graphs) and respective strengths and drawbacks.
After an analysis of the literature we observed that many tasks are under-explored or addressed without a systematic evaluation. So, we introduced a formal definition for the task of semantic change characterization, which is a more complex instance of the semantic change identification problem. This new task is of extreme relevance for the field of LSC as this study and other surveys defended.
The suggested tasks are crucial for both linguistic change research and NLP; numerous tasks, such as sentiment analysis (orientation), semantic similarity (relation), and word sense inference (dimension), call for a deeper understanding of these characteristics of representation. Additionally, improved knowledge and interpretability of these systems result from an awareness of how changes affect the representation of words. Specially for word embeddings, better understanding how this changes affect the geometries could improve our knowledge in the semantics of vectors.
As a future work we expect to investigate the relation pole to better differentiate the categories of change, and further explore the four word representational methods, how they affect and are related to to the three poles of change. Another important improvement is a consolidation of datasets and metrics where semantic change characterization could be compared across methodologies.
Finally, the emergence of new-generation language models is acknowledged as opening up new possibilities for identifying and characterizing semantic changes, though this area of research is still under-explored. The conclusion suggests avenues for future research, addressing the identified challenges and leveraging the capabilities of advanced language models.
Acknowledgments
The research reported in this publication was supported by the Luxembourg National Research Fund (FNR), project D4H grant number PRIDE21/16758026